OpenStudy (anonymous):

What is the solution to the equation 256 = (1/2x -3)3

OpenStudy (anonymous):

help

OpenStudy (anonymous):

256=(1/2x-3)^3

OpenStudy (anonymous):

The lenght of the diagonal of a square is 8 inches. What is the area of the square?

mathslover (mathslover):

Which question do you want help for?

OpenStudy (anonymous):

both and you can start with anyone

mathslover (mathslover):

Okay, first one will be discussed first. So, we have : \(256 = (\dfrac{1}{2x} -3)^3\) , or is x in numerator?

OpenStudy (anonymous):

thats right

OpenStudy (anonymous):

hello

mathslover (mathslover):

Sorry, have to go for breakfast. Can you wait fro 5-6 minutes?

OpenStudy (anonymous):

yes

mathslover (mathslover):

Okay to solve for this question, take cube root both sides.

mathslover (mathslover):

What is cube root of 256?

OpenStudy (anonymous):

16

mathslover (mathslover):

No.. Its not 16 as 16^3 is not 256. It is 6

mathslover (mathslover):

so, we get 6 = \(\dfrac{1}{2x} -3 \)

OpenStudy (anonymous):

x = -26 only x = 38 only x = -26 and x = 38 x = 13 and x = 19